A bus started from rest and moved with uniform acceleration. It acquired a speed of 60 m/s after 100 seconds. What is the acceleration and the distance traveled after 100 seconds?


1
Expert's answer
2020-11-11T08:55:58-0500

We can find the acceleration of the bus from the kinematic equation:


"v=v_0+at,""a=\dfrac{v-v_0}{t}=\dfrac{60\ \dfrac{m}{s}-0\ \dfrac{m}{s}}{100\ s}=0.6\ \dfrac{m}{s^2}."

We can find the distance traveled after 100 seconds from another kinematic equation:


"d=v_0t+\dfrac{1}{2}at^2,""d=0\ \dfrac{m}{s}\cdot 100\ s+\dfrac{1}{2}\cdot 0.6\ \dfrac{m}{s^2}\cdot (100\ s)^2=3000\ m."

Answer:

"a=0.6\ \dfrac{m}{s^2}."

"d=3000\ m."